1. The number of bank robberies that occur in Lagos is modeled to be Poisson distributed with mean of 1.8 per day. Find the probability of getting exactly three robberies in a day.

a. 0.3213

b. 0.1607

c. 0.8393

d. 0.4137



1
Expert's answer
2021-01-21T00:19:26-0500

In the poison distr ibution a discrete random variable "X" has the following probability mass function,

"P(X=x)=\frac{e^{-\lambda }.\lambda ^x}{x!}" , where "\lambda" is the mean of the distribution and "x=0,1,2,3.....\infty"

Given that "\lambda =1.8"

Hence the required probability, "P(X=3)=\frac{e^{-1.8 }.(1.8)^3}{3!}=0.1607"

So probability of getting exactly three robberies in a day is "=0.1607"